integration of √cot x

Grade:12

1 Answers

Anoopam Mishra
8 years ago
\int\sqrt{cot(x)} = \frac{1}{2\sqrt{2}}(-log(cot(x) - \sqrt{2}\sqrt{cot (x)} +1) + log(cot(x) + \sqrt{2}\sqrt{cot(x)} + 1) + 2 tan^{-1}(1-\sqrt{2}\sqrt{cot(x)}) - 2tan^{-1}(\sqrt{2}\sqrt{cot(x)} + 1) + constant